OpenStudy (anonymous):

Which of the following represents an arithmetic sequence? A 5, -10, 20, -40 B 3, 1, 4, 2 C 2, 4, 8, 16 D 3, 5, 7, 9,

OpenStudy (campbell_st):

arithmetic sequences keep changing by the same amount... called the common difference. normally found by checking the terms in the sequence. \[T _{2} - T _{1} = T _{3} - T _{2} = T _{4} - T _{3}=....,.\]

OpenStudy (anonymous):

sooo..

OpenStudy (anonymous):

how do i apply tht to the question

OpenStudy (campbell_st):

so in A -10 - 5 = -15 20 - (-10) = 30 different values not arithmetic B 1 -3= -2 4 - 1 = 3 not arithmetic C 4 - 2 = 2 8 - 4 = 4 not arithmetic D 5 -3 = 2 7 - 5 = 2 9 - 7 = 2 the common difference is 2... so the sequence is arithmetic

OpenStudy (anonymous):

thank you! :)